OpenStudy (anonymous):

Part a.) Find the slope of the line that contains the points, (1, 2) and (2, 5). Part b.) Using the slope that you found in part a and one of the given points, write the equation of the line. Please write your final answer in slope-intercept form.

OpenStudy (anonymous):

start with point-slope form: y - y1 = m(x - x1) then you need to "plug" in your points 2 - 5 = m(1 - 2) now solve for m which represents your slope -3 = m(-1) now divide both sides by -1 3 = m now that you know the slope we can start on part b slope intercept form is: y = mx + b we need to use one of our points to find b 2 = 3(1) + b now solve for b which represents our y-intercept 2 = 3 + b subtract the 3 from both sides: -1 = b now we can plug in b and get our equation in slope intercept form y = 3x -1